OpenStudy (anonymous):

Determine whether f(x) and g(x) are inverses of each other: f(x) =2/3 x+6 g(x) = 3/2 x-9

OpenStudy (amistre64):

if you use one inside the other and it = x; they are inverse, but to be inverse both ways they have to both be able to nest in each other and get an x

OpenStudy (amistre64):

f(g(x)) = 2/3(3/2 x-9)+6 = x -6+6 = x g(f(x)) = 3/2 (2/3 x+6)-9 = x +9-9 = x

OpenStudy (amistre64):

they are inverses of each other them